La Mode Design Inc., a high-fashion women's dress manufacturer, is planning to market a new cocktail dress for the coming season. La Mode Design supplies retailers in Toronto, Montreal, and the Atlantic provinces.

Four metres of material are laid out for the dress pattern. After cutting, some material remains, which can be sold as remnants. The company could also use the leftover material to manufacture a matching cape and handbag. However, if it uses the leftover material for the cape and handbag, more care will be needed in the cutting, and the cutting costs will therefore increase.

The company expects to sell 1,250 dresses if a matching cape and handbag are not available. La Mode Design's market research reveals, however, that dress sales would be 20% higher if a matching cape and handbag were available. The market research indicates that the cape and/or handbag could not be sold individually but only as accessories with the dress. The various combinations of dresses, capes, and handbags that retailers will sell are as follows:

Complete sets of dress, cape, and handbag

70%

Dress and cape

6%

Dress and handbag

15%

Dress only

9%

Total

100%

The material used in the dress costs $12.50 a metre, or $50.00 for each dress. The cost of cutting the dress if the cape and handbag are not manufactured is estimated at $20.00 a dress, and the resulting remnants can be sold for $5.00 for each dress cut out. If the cape and handbag are to be manufactured, the cutting costs will be increased by $9.00 per dress. There will be no saleable remnants if the capes and handbags are manufactured in the quantities estimated.

The selling prices and the costs to complete the three items once they are cut are as follows:

Selling Price per Unit

Unit Cost to Complete*

Dress

$200.00

$80.00

Cape

    27.50

  19.50

Handbag

      9.50

    6.50

Instructions

a.  Prepare La Mode Design's incremental analysis for manufacturing the capes and handbags with the dresses.

b.  Based on your analysis, what decision should management make?

c.  Identify any qualitative factors that could influence the company's decision to manufacture the capes and handbags that match the dresses.
Answer #1

Answer to (a)

Summary of Question -

Normal Production = 1250 dresses

Material required for each dress = 4m | Cost of Cloth = $ 12.5 per m | Cost of cloth per Dress = $ 50

Cutting Cost per Dress = $20

Incremental Analysis -

Calculation of Incremental Revenue Amount
A. Revenue if Caps and Handbags are also manufactured (WN 1) $ 343,490.5
B. Revenue when Caps and Handbags are NOT manufactured (WN 2) $ 256,250
C. Incremental Revenue (A-B) $ 87,242.5
D Incremental Cost (WN 3) $ 81,517.5
Profit on Manufacturing Cap and Handbag (C- D) $ 5725

WN 1

Particular No of Units Price per unit Amount
Complete set of Dress, Cap, Handbag 1500*70% = 1050 200+27.5+9.50 =237 248880
Dress and Cap 1500*6% = 90 200+27.5 = 227.5 20475
Dress and Handbag 1500*15% = 225 200+9.5 = 209.5 47137.5
Dress Only 1500*9% = 135 200 27000

WN 2

Revenue from Sale of Dress (1250*200) = 250000

Revenue from Sale of Remnants (1250*5) = 6250

Total Revenue (250000+6250) = 256250

Calculation of Incremental Cost (WN 3) -

Particular Calculation Amount Description
Incremental Cloth Cost (1500-1250)*50 12500 20% increase in Sale of Dress will require more cloth
Incremental Cutting Cost 1500* $ 9 per dress + 250 * 20 per dress 18500 Cutting cost is required on all units as there will be no saleble value of remenats if caps and handbags are manf. Which means all the cloth will be utilised
Incremental Unit cost of Dress 250*80 20000 20% increase in Sale of Dress will require more Unit cost
Incremental Unit cost of Cap 1140*19.5 22230 Since Cap will be sold 1140 units only ( 1050 with complete set and 90 with Dress)
Incremental Unit cost of Handbag 1275*6.5 8287.5 Since Handbag will be sold 1275 units only ( 1050 with complete set and 225 with Dress)
Total 81517.5

Answer to b

Base on the Incremental Analysis Incremental Revenue is More then Incremental cost and therefore management should go for manufacture of Caps and Handbags as it will result into Profit of $ 5725.

Answer to c

Qualitative factors that influence company's decision to manufacture matching Caps and Dresses are -

1. As Market research provides that with matching Caps and Handbags Demand for Dress will go up by 20%

2. Efficient use of resources and less wastage which will help co in environmental grounds

3. Help in diversification of company's products